OpenStudy (anonymous):

which method would be the simplest way to solve the system please help me? i need the awnser y=1/2x 2x+3y=28 a.graphing b.substituting. c.eliminating d.distributive

OpenStudy (nincompoop):

distributive will be the simplest way 2x + 3*(1/2 x) = 28

OpenStudy (nincompoop):

I mean substitute

OpenStudy (anonymous):

are you sure

OpenStudy (nincompoop):

VERY SURE like 10000000000000% sure

OpenStudy (anonymous):

thanks :3

OpenStudy (nincompoop):

I showed you the actual method

OpenStudy (nincompoop):

this is substitution 2x + 3*(1/2 x) = 28 where y = 1/2 x

OpenStudy (nincompoop):

thanks for the medal, I don't need it :P